### v2

# MANGO: A Benchmark for Evaluating Mapping and Navigation Abilities of Large Language Models

Abstract
Large language models such as ChatGPT and GPT-4 have recently achieved astonishing performance on a variety of natural language processing tasks. In this paper, we propose MANGO, a benchmark to evaluate their ability to perform text-based mapping and navigation. Our benchmark includes $53$ mazes taken from a suite of textgames: each maze is paired with a walkthrough that visits every location but does not cover all possible paths. The task is question-answering: for each maze, a large language model reads the walkthrough and answers hundreds of mapping and navigation questions such as “How should you go to Attic from West of House?” and “Where are we if we go north and east from Cellar?”. Although these questions are easy for humans, it turns out that even GPT-4, the best-to-date language model, performs poorly when answering them. Further, our experiments suggest that a strong mapping and navigation ability would benefit the performance of large language models on relevant downstream tasks, such as playing textgames. Our MANGO benchmark will facilitate future research on methods that improve the mapping and navigation capabilities of LLMs. We host our leaderboard, data, code, and evaluation program at https://mango.ttic.edu and https://github.com/oaklight/mango/.
1 Introduction
Mapping and navigation are fundamental abilities of human intelligence (Spiers & Maguire, 2006; Epstein et al., 2017). Humans are able to construct maps—in their minds (Epstein et al., 2017) or on physical media like paper—as they explore unknown environments. Following these maps, humans can navigate through complex environments (Spiers & Maguire, 2006; Spiers & Gilbert, 2015; Javadi et al., 2017), making informed decisions, and interact with their surroundings. Such abilities empower humans to explore, adapt, and thrive in diverse environments. An example is remote (e.g., deep-sea) exploration for which humans have drawn upon their intuition to develop algorithms that enable robots to autonomously navigate and map their surroundings based only on onboard sensing.
Do large language models (LLMs) possess such abilities? In this paper, we investigate this research question by creating a benchmark and evaluating several widely used LLMs. Our MANGO benchmark is the first to measure the mapping and navigation abilities of LLMs. It includes $53$ complex mazes, such as the one visualized in Figure 1. It pairs each maze with hundreds of destination-finding questions (e.g., “Where will you be if you go north, north, and then up from Altar?”) and route-finding questions (e.g., “How do you reach Dome Room from Altar?”). For each maze, the language model has to answer these questions after reading a walkthrough of the maze. Many questions involve possible routes that are not traced during the walkthrough, making the benchmark challenging. In our experiments, GPT-4 only correctly answered half of the route-finding questions, performing disastrously on the difficult questions (e.g., those involving long and unseen routes). MANGO will facilitate future research in improving the mapping and navigation abilities of LLMs.
Figure 1: Map of Zork-I. Arrows denote the direction of travel during the walkthrough, while the reverse direction is unseen but may be possible. Note that it is a 3D map projected onto a 2D plane so up may not point upward in the 2D visualization (e.g., Rocky Ledge to Canyon View).
Another contribution of MANGO is to draw a novel connection between natural language processing and robotics. There has been significant interest in employing LLMs to endow intelligent agents (including robots) with complex reasoning (Yang et al., 2023). Aligning with this interest, MANGO enables the investigation of the LLMs’ capabilities in simultaneous localization and mapping (SLAM) within text-based worlds. Focusing on this aspect, our work stands out and complements previous SLAM-related research, which predominantly relies on richer sensory inputs (e.g., vision and LiDAR).
2 MANGO: A Benchmark for Text-Based Mapping and Navigation
Our MANGO benchmark measures the mapping and navigation capabilities of LLMs. It leverages a suite of text-based adventure games that offer expert-designed complex environments but only require simple actions. Figure 1 is an example: it was drawn according to the first 70 steps of the walkthrough of Zork-I, which can be found in LABEL:lst:walkthrough. This map is imperfect: the annotator had to draw the only Kitchen twice to avoid a cluttered visualization; the Living Room was incorrectly placed outside the House. However, equipped with this map, one could correctly answer questions about any route in the maze such as “How do you reach Dome Room from Altar?” and “Where will you be if you go north, north, and up from Altar?”. The walkthrough has not traced a route from Altar to Dome Room, but humans possess the remarkable capacity to plan a route by identifying the three individual steps—which the walkthrough has covered—from Dome Room to Altar and retracing those steps. MANGO tests whether a large language model can perform the same kind of reasoning. Particularly, when evaluating a language model, we first let it read a walkthrough like LABEL:lst:walkthrough and then ask it questions like those in LABEL:lst:dfq and LABEL:lst:rfq. A question like LABEL:lst:dfq is a destination-finding (DF) question, and a question like LABEL:lst:rfq is a route-finding (RF) question. Users of MANGO have the flexibility to phrase the DF and RF questions in their own ways: as shown in LABEL:lst:dfqraw and LABEL:lst:rfqraw, we provide the skeletons of these questions, which users can plug into their own templates.
STEP NUM: 0
ACT: Init
OBSERVATION: West of House
You are standing in an open field west of a white house, with a boarded front door. There is a small mailbox here.
STEP NUM: 1
ACT: south
OBSERVATION: South of House
You are facing the south side of a white house.
STEP NUM: 2
ACT: east
OBSERVATION: Behind House
You are behind the white house. A path leads into the forest to the east. In one corner of the house there is a small window which is slightly ajar.
$\vdots$
STEP NUM: 70
ACT: east
OBSERVATION: Gallery
This is an art gallery. Most of the paintings have been stolen by vandals with exceptional taste. The vandals left through either the north or west exits. Fortunately, there is still one chance for you to be a vandal, for on the far wall is a painting of unparalleled beauty.
Example 1: An example of Zork-I walkthrough.
Starting from Altar, perform actions [north, north, up], where are you now?
Example 2: A destination-finding question.
How can you go from Altar to Dome Room?
Example 3: A route-finding question.
S: Altar # starting location
A: north, north, up # list of actions
Example 4: Skeleton of DF question in LABEL:lst:dfq.
S: Altar # starting location
D: Dome Roomm # destination
Example 5: Skeleton of RF question in LABEL:lst:rfq.
S: Altar
A: north
D: Temple
S: Temple
A: north
D: Torch Room
S: Torch Room
A: up
D: Dome Room
Example 6: Full route of LABEL:lst:dfq and LABEL:lst:dfqraw.
author=DP,color=blue!20,size=,fancyline,caption=,]in rebuttal, we promised a complete example of walkthrough, we may use partial zork1 as appendix or some other ones with less step
2.1 Maze Collection: From Game Walkthroughs to Mazes
Our mazes are taken from the textgames in the Jericho game suite (Hausknecht et al., 2020). The main release of Jericho includes $57$ popular textgames as well as a program that can generate walkthroughs for 56 of them. The original walkthrough of a game is a list of actions (such as east, north, and open door) that one could execute to efficiently complete the game. We enhanced each walkthrough by executing the sequence of actions and augmenting each step with the new observation (i.e., the text feedback that the game engine provides after the action is executed). Unless explicitly specified, the word “walkthrough” refers to the enhanced, but not original, walkthroughs (such as LABEL:lst:walkthrough) throughout the paper. More details about walkthroughs can be found in section A.1.
In a walkthrough, not every action triggers a location change: it may update the inventory (such as take lamp and drop pen) or time (such as wait). For each game, we read the walkthrough, labeled the actions (such as east and up) that change the locations, and made note of the names of the locations (such as Temple and Altar). This annotation is nontrivial and can not be automated. We had to pay extra attention to appropriately handle the tricky cases including: ① the name of a location may be mentioned in a rich, but distracting context (e.g., the context may have ten paragraphs and hundreds of words with the name briefly mentioned in the middle); ② a location may be visited multiple times, so we need to assign the same name to all its mentions; ③ different locations may be referred to with the same name in the textual feedback, so we need to rename them in a sensible way.
The location name resolution (see section A.2 for a full procedure) results in a maze for each game. Three of the games have no location change, and so we left them out, resulting in $53$ mazes. We store each maze as a directed graph: each node is a named location (e.g., Altar); each directed edge is a movement (e.g., north); and each node-edge-node combination is a location-changing step that was followed in the walkthrough. Note that a graph may be cyclic since the walkthrough may trace back-and-forth between locations (e.g., Temple and Egyptian Room in Figure 1).
2.2 Generation of Question Skeletons: Traversing Mazes and Imputing Edges
To generate DF and RF skeletons for a maze, a naive approach is to perform brute-force traversal. First, we collect all the possible S-P-D tuples, where S and D are locations and P is a simple path from S to D. A simple path is a directed path that does not visit any location more than once. This “simple” restriction ensures that we will have a finite number of S-P-D tuples. LABEL:lst:route is a simple path of 3 S-A-D edges from Altar to Dome Room. Each unique S-P-D tuple gives a unique DF skeleton: e.g., LABEL:lst:dfqraw is obtained from LABEL:lst:route. Each unique S-P-D tuple gives an RF skeleton as well, such as LABEL:lst:rfqraw obtained from LABEL:lst:route. However, the same RF skeleton may be obtained from other tuples since there may be multiple possible simple paths between the same pair of locations S and D. As a consequence, we may end up with fewer RF questions than DF questions for a given maze.
The particular DF and RF questions in LABEL:lst:dfq and LABEL:lst:rfq are challenging to large language models, since they involve actions—such as going north from Altar to Temple—that are not covered in the walkthrough. Answering such hard questions requires a deeper understanding of the spatial relationships between locations. However, also because these steps are not in the walkthrough, the skeletons in LABEL:lst:dfqraw and LABEL:lst:rfqraw can not be obtained through a naive traversal of the directed graph in Figure 1. That is, we have to traverse an extended graph that includes imputed edges. An imputed edge denotes a valid step that is not explicitly mentioned in the walkthrough, such as going north from Altar to Temple (i.e., Altar-north-Temple). Most mentioned edges involve directional moves (e.g., up, east), so reversing them is a straightforward way to impute new edges. We manually examined other edges: for some of them, we proposed intuitive reverses (such as exit for enter); for the others (e.g., pray), no reverse could be found. We then examined the imputed edges through real game play and discarded those failing to cause the expected location changes. section A.3 documents the full procedure of edge imputation and examination.
After extending all the mazes in our benchmark, we collected 21046 DF skeletons and 14698 RF skeletons by traversing the extended graphs. Being evaluated on a maze, the LLM may not be able to consume the entire walkthrough in its context window. That is, we may only feed it an appropriate prefix of the walkthrough (e.g., the first $70$ steps for Zork-I as shown in LABEL:lst:walkthrough), leaving some of the DF and RF skeletons unanswerable given that prefix. Therefore, our benchmark provides the ANSWERABLE label (an integer) for each skeleton such that this skeleton is only answerable if the maximum STEP NUM in that prefix (e.g., $70$ in LABEL:lst:walkthrough) is greater than or equal to its ANSWERABLE label. Furthermore, given a walkthrough prefix, an answerable skeleton may be easy or hard, depending on whether it involves edges that are not covered in the prefix. Precisely, a DF skeleton is considered to be easy if all the S-A-D edges in its corresponding simple path are covered in the walkthrough prefix; an RF skeleton is easy if the shortest simple path from its starting location to its destination only involves the S-A-D steps covered in the prefix. When a longer walkthrough prefix is used, more answerable questions tend to become easy. Our benchmark provides the EASY label (also an integer) for each skeleton: a skeleton is easy if the maximum STEP NUM in the walkthrough prefix is no smaller than its EASY label; otherwise, it is a hard skeleton. Table 3 in Appendix A documents the statistics of the full dataset, such as the number of locations and the number of skeletons. Tables 5, 8, 6 and 7 in Appendix B shows the information about the data on which each LLM was evaluated in our experiments.
2.3 Evaluation Program
The evaluation program in our benchmark implements a range of evaluation and analysis methods. Reading the model-generated answers, it can return a set of evaluation scores together with rich analysis. In this section, we introduce the most important scores used in our main experiments. Other scores are discussed in section A.6, with their related experiments presented in Appendix C.
For DF questions, the most straightforward evaluation is the success rate: i.e., the fraction of questions that the language model answers correctly. What answers will be considered to be correct? A strict criteria is that the model answer is correct if and only if it exactly matches the ground-truth location name. However, due to the variability of natural language, a correct answer may not exactly match the ground-truth. For example, the model may yield The House or That House when the ground-truth location name is just House. To account for such cases, we generalize the success rate to allow partial matches. Given a model answer $\hat{A}$ and the ground-truth answer $A$, we compute their (character-level) edit-distance $d$ and define a correctness score $c\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}1-d/\ell$ where $\ell$ is the length of the longer answer. The score is $\in[0,1]$: when the answer exactly matches the ground-truth, we have $c=1$; if they have no character overlap at all, then $c=0$. We then define the success rate to be the sum of the correctness scores over all the questions, divided by the number of questions.
For RF questions, the main metric is still the success rate, but the definition of “success” is different from that for DF questions. Note that an answer to an RF question is a sequence of moves. We consider an answer to be correct if and only if it can reach the destination after our evaluation program executes it in the maze. A correct answer to an RF question may not be a good path: it doesn’t have to be the shortest; it doesn’t even have to be a simple path. It is possible that an LLM-generated move is meaningful but doesn’t exactly match any valid move in the graph: e.g., the LLM may give walk south, which means the same as south. Therefore, when executing a model-generated move, our evaluation program will select the closest (i.e., smallest edit-distance) valid move.
3 Experiments
In this section, we present the results our evaluation of several widely used LLMs.
3.1 Experiment Setup
The evaluated models are: GPT-3.5-turbo (Brown et al., 2020; Stiennon et al., 2020; Gao et al., 2022), GPT-4 (OpenAI, 2023), Claude-instant-1 (Anthopic, 2023a), Claude-2 (Anthopic, 2023b), Llama-2 with 13B parameters (Touvron et al., 2023b), and RWKV with 14B parameters (Peng et al., 2023). For GPTs and Claudes, we used the prompt templates in LABEL:lst:maptemplate and LABEL:lst:navtemplate, converting the DF and RF skeletons like LABEL:lst:dfqraw and LABEL:lst:rfqraw into LLM-friendly questions like LABEL:lst:dfq and LABEL:lst:rfq. The templates were carefully designed and examined through pilot experiments, in order to ensure that we do not underestimate the models on our benchmark. In our templates, each question starts with a list of legal actions, followed by a list of reachable locations; these lists help mitigate the hallucination of language models. The templates ask the model to spell out the entire trajectory including all the intermediate locations. This design is inspired by Chain-of-Thought prompting (Wei et al., 2022): eliciting an LLM to give its entire reasoning process tends to improve its overall performance on downstream tasks. In addition, it allows us to conduct a deeper evaluation and analysis, such as the reasoning accuracies of the models (see sections A.6 and C). Note that our templates request the model to form its answer as a list of Python dictionaries with specific key names. We found that this restriction encourages the model to generate structured answers—which are easy to parse and analyze—as well as improves its performance. For Llama-2 and RWKV, we made moderate revisions to the prompts in order to generate well-structured answers as well as optimize for their performance.
The allowed actions are:
The list of places are:
Starting from S, perform a list of
actions [A], where are you now?
Describe the trajectory in a Python list of Python dictionaries with keys prev_node’, node and action’.
Start your response with ’[’.
Example 7: Our DF template.
The allowed actions are:
The list of places are:
How can you go from S to D?
Describe the trajectory in a Python list of Python dictionaries with keys prev_node’, node and action’.
Start your response with ’[’.
Example 8: Our RF template.
For GPT-3.5, we experimented with the 4K version, which can consume 4096 tokens in its context window. This context limit restricts the length of the walkthrough that it can read, and the number of DF and RF questions that it can answer. Table 5 shows the statistics about the walkthrough prefix and questions that GPT-3.5 used for each maze. For GPT-4, Claude-1 and Claude-2, we used the same walkthrough prefixes and questions as GPT-3.5 for a fair comparison. we used the same walkthrough prefixes and questions as GPT-3.5 for a fair comparison. Llama-2 has a $4096$ context window as well. But its tokenizer is different from GPTs’ so we evaluated it on a slightly different set of questions. RWKV is capable of handling infinite context. For each maze, we experimented it with the $70$-step prefix of the walkthrough so that its set of answerable questions includes all the questions answered by all the other models. We also evaluated Llama-2 and RWKV in a simplified setting, where the observation at each step of the walkthrough only includes the location name but nothing else. For example, at STEP 1 of the simplified LABEL:lst:walkthrough, OBSERVATION only has South of House and everything else (i.e., Your are...) is omitted. We refer to Llama-2 and RWKV with the simplified walkthroughs as Llama-2-S and RWKV-S, respectively. More details about the experiment setup are in Appendix B.
3.2 Main Results
Figure 2 presents the success rates of all models. For each kind of question (i.e., DF or RF), we show the results on easy and hard questions separately. As we can see, GPT-4 significantly outperforms all the other models on all kinds of questions. However, it only correctly answers half of the RF questions, far worse than what a human could do: in our experiments, humans perfectly answered a randomly sampled set of questions.
(a) On easy (left) and hard (right) DF questions.
(b) On easy (left) and hard (right) RF questions.
Figure 2: Success rates of the examined models on (2(a)) DF and (2(b)) RF questions, averaged over all $53$ mazes. Appendix C provides similar graphs (e.g., Figure 6) for other evaluation metrics.
Method RWKV Llama-2 Claude-1 Claude-2 GPT-3.5 GPT-4 $\overline{{\rm HARD}}|$
RWKV * 0.20 | 0.24 0.19 | 0.41 0.19 | 0.51 0.19 | 0.33 0.19 | 0.62 *
Llama-2 0.43 | 0.20 * 0.24 | 0.41 0.24 | 0.45 0.24 | 0.31 0.24 | 0.66 *
Claude-1 0.74 | 0.19 0.78 | 0.41 * 0.36 | 0.44 0.38 | 0.32 0.36 | 0.57 *
Claude-2 0.82 | 0.19 0.85 | 0.41 0.81 | 0.72 * 0.44 | 0.32 0.44 | 0.58 *
GPT-3.5 0.59 | 0.19 0.61 | 0.42 0.57 | 0.74 0.57 | 0.83 * 0.32 | 0.59 *
GPT-4 0.86 | 0.19 0.90 | 0.42 0.84 | 0.72 0.83 | 0.81 0.86 | 0.57 * *
|$\underline{{\rm EASY}}$
(a) Pairwise comparison on easy (lower left) and hard (higher right) DF questions.
Method RWKV LlaMa-2 Claude-1 Claude-2 GPT-3.5 GPT-4 $\overline{{\rm HARD}}|$
RWKV * 0.00 | 0.00 0.00 | 0.13 0.00 | 0.20 0.00 | 0.03 0.00 | 0.54 *
Llama-2 0.02 | 0.02 * 0.00 | 0.16 0.00 | 0.21 0.00 | 0.05 0.00 | 0.46 *
Claude-1 0.36 | 0.01 0.34 | 0.03 * 0.11 | 0.19 0.13 | 0.03 0.11 | 0.45 *
Claude-2 0.49 | 0.01 0.46 | 0.03 0.47 | 0.33 * 0.20 | 0.03 0.19 | 0.46 *
GPT-3.5 0.16 | 0.01 0.17 | 0.03 0.15 | 0.36 0.15 | 0.50 * 0.03 | 0.48 *
GPT-4 0.57 | 0.01 0.56 | 0.03 0.55 | 0.33 0.55 | 0.47 0.58 | 0.15 * *
|$\underline{{\rm EASY}}$
(b) Pairwise comparison on easy (lower left) and hard (higher right) RF questions.
Table 1: Success rates on DF and RF questions broken down into pairwise comparisons. In each table, the cell of row-A and col-B contains the success rates of the models—in the format of A | B—on the intersection of the questions that A and B answered individually. The lower left triangle displays the results on easy questions, while the upper right triangle shows the results on hard questions.
Note that each model was evaluated on its specific set of questions determined by the length and format of the walkthrough it read. To be fair, we also compared each pair of models on the intersection of the questions that they answered. The results are presented in Table 1: as we can see, GPT-4 and GPT-3.5 consistently outperform the other models and GPT-4 significantly outperforms GPT-3.5.
More results are in Appendix C, including results on other evaluation metrics (e.g., weighted success rates) and comparison between Llama-2 with Llama-1 and Llama-2-chat. We also explored an alternative approach that first maps a walkthrough to a symbolic graph and then uses a search algorithm to answer the given (DF or RF) question. Results and analysis of this approach can be found in section C.4. Overall, we found it to be very challenging to translate natural language walkthroughs into searchable symbolic graphs in the first place, making this approach not promising.
3.3 Analysis of GPTs
Now, we focus our analysis on the best model, namely GPT-4. Particularly, we would like to understand the improvements of GPT-4 over GPT-3.5 as well as its current bottlenecks, shedding light on opportunities for future improvements.
By analyzing the errors of GPT-3.5 and GPT-4, we discovered that these models occasionally hallucinate nonexistent locations or edges. Once they made such a mistake at any step of reasoning, they would be misled and deviate from the correct path towards the correct answer. Furthermore, we found that the mazes are not equally difficult for the models. Figure 3 displays the success rates of the GPT models broken down into their per-game results. In Figure 3, each dot is a maze: the $x$-axis coefficient is the performance of GPT-3.5 on this maze while the $y$-axis is that of GPT-4.
(a) Easy DF.
(b) Hard DF.
(c) Easy RF.
(d) Hard RF.
Figure 3: Success rates of GPT-3.5 and GPT-4 broken down into individual games. Figure 7 in Appendix B provides a similar visualization of reasoning accuracy.
As we can see, the success rates of the models vary across different mazes as well as across different kinds of questions. GPT-4 consistently outperforms GPT-3.5 across nearly all the mazes. The only exception is Seastalker: there are too few hard DF questions for this maze, and thus it is a noisy outlier. Apparently, both GPTs tend to work better on easy questions than on hard questions. However, some mazes seem to be particularly challenging to GPT-4, such as Zenon and OMNIQuest.
What makes those mazes challenging?
We collected some important statistics about the mazes and analyzed their correlation with the success rates of the models. To understand the success rates on the easy questions, it is interesting to investigate:
  • number of locations (# locations) and number of explicit edges (# exp edges). They directly measure the size of a maze, which may be a key indicator of its difficulty.
  • number of potentially confusingly named locations (# conf locations). Recall from section 2.1 that different locations may have similar or related names, which may confuse a language model. To quantify the number of confusingly named locations, we compute a confusion score for each location, and then sum the scores across all the locations. For a location name $A$, the confusion score is defined to be the maximum word-level edit distance between $A$ and any other location name in the maze, divided by the maximum word-level length of the pair of location names being compared. Technically, it is $\max_{B}\left(\text{edit-distance}(A,B)/\max(\text{len}(A),\text{len}(B))\right)$, and it is $\in[0,1]$.
  • average length of the easy simple paths (avg len easy), i.e., the simple paths that do not include any imputed edges. A longer path may tend to be more difficult for models.
  • average number of words in the scene descriptions (avg len scene). The walkthroughs exhibit very diverse styles: for some of them, the text description for each scene is very concise and the name of each location is appropriately highlighted; for others, each description may be verbose (e.g., ten paragraphs and hundreds of words) and the location names are often not obvious from the contexts. It is useful to analyze whether a long scene description poses a challenge for the models.
In order to understand the models’ performance on hard questions, we analyze the effects of the variables above (except avg len easy) as well as the following:
  • number of imputed edges (# imp edges);
  • average length of hard—i.e., involving imputed edges—simple routes (avg len hard);
  • average number of imputed edges in the hard simple routes (avg # imp in hard).
We use regression analysis to understand the effects of these variables on model performance. In particular, for each model on each type of question (DF or RF, easy or hard), we ran single-variable linear regression to understand how the success rate varies with each variable. Detailed results (e.g., coefficients and $p$-values) are section C.3. Overall,
  • on easy questions, GPTs are significantly influenced by the size of the maze, the confusion level of location name, and the path length. The $p$-values are extremely small.
  • on easy questions, the average length of the scene descriptions does not have a significant effect on the performance of GPT-3.5, but interestingly has a significant positive effect on GPT-4’s performance. It is perhaps because GPT-4 possesses a strong capability to understand texts and can leverage the rich contexts in each description. This allows it to better distinguish confusingly named locations and establish a better internal representation of the map. However, this richness seems to confuse GPT-3.5 and impede its ability to create a good internal representation of the maze, possibly due to GPT-3.5’s weaker overall language understanding capabilities.
  • on hard questions, the variables do not significantly affect the performance of GPT-3.5. Note that GPT-3.5 yields very low success rates when answering the hard DF and RF questions. GPT-3.5 seems to struggle when it has to reason about a path with any number of imputed edges, making the effect of other factors less important to its performance.
  • on hard questions, GPT-4 exhibits a stronger ability to handle paths with imputed edges, compared to GPT-3.5. However, it will experience difficulties when the challenge of inferring imputed edges is amplified by other factors such as the size of the maze or the length of the path. As a result, nearly all the variables have significant effects on GPT-4.
The results of our regression analysis are consistent with the plots in Figure 3. For example, both Zenon and OMNIQuest stay at the lower-left corners of the hard-question plots in Figure 3 since their mazes are particularly challenging to both GPT-3.5 and GPT-4: they both have substantially larger numbers of imputed edges than the other mazes; OMNIQuest also has more locations. Wishbringer and Lost Pig have several imputed edges, but their paths are short, so they fall in the upper-left corners of the hard-question plots in Figure 3.
3.4 Human Performance
We measured human performance on a subset of our data. This subset includes 30 DF questions (21 easy and 9 hard) and 31 RF questions (20 easy, 11 hard). The student authors participated in two rounds of evaluation: in the first round, each author answered a random split of the questions, and all the questions were answered; in the second round, we randomly sampled 10 DF and 10 RF questions and let each of them be re-evaluated by a different author. The second round allows us to analyze human agreement; this analysis can be found in section C.6.
Task Difficulty Success Rate Reasoning Accuracy
Route Finding All (31) 0.8211 0.6129
Easy (20) 0.7727 0.5500
Hard (11) 0.9091 0.7273
Destination Finding All (30) 1.0000 0.5667
Easy (21) 1.0000 0.6667
Hard (9) 1.0000 0.3333
Table 2: Human Evaluation Results
As shown in Table 2, humans generally achieve high success rates on both RF and DF questions, and exhibit small to no difference across easy and hard questions. This is interesting since LLMs tend to struggle on hard questions. Reasoning accuracies are lower than success rates; this is because human answers exhibit a relatively high variability in writing (e.g., location names with special symbols).
Despite the high scores, answering the questions is actually not trivial for human raters. Answering each DF question takes an average of 15 minutes, while answering each RF question takes an average of 30 minutes. All the human raters needed to take notes (e.g., drawing maps) while completing these tasks.
3.5 Does Mapping and Navigation Ability Matter in Downstream Tasks?
Now we present a case study showing that a strong mapping and navigation ability of an LLM would benefit it in downstream tasks. In particular, we selected 284 minigames in the Jericho game suite, and investigated how the map knowledge may improve the performance of an LLM in playing these minigames. Each minigame is a selected prefix of a walkthrough from one of 53 textgames; the selection criterion is that the best action to take at this step is a movement. In other words, each minigame is a scenario where the LLM has to figure out the best action to take given its previous actions and observations (i.e., the prefix of walkthrough up to the current step). This task is different and more challenging than answering the DF and RF questions: the LLM is not explicitly given a route (as in DF questions) or a destination (as in RF questions), but has to spontaneously figure out which action may contribute to its long-term goal.
# previous actions and observations
Small local map info: if you want to go to North of House, you should go south; if you want to go to Up a Tree, you should go up; if you want to go to Altar, you should go west.
Consider what you should do next, and choose one appropriate action from the valid actions list: [up, take on egg, put down egg, go around forest, throw egg at tree, open egg with all, north, south, west, east]
Please just tell me the selected action without any extra words.
Example 9: A prompt of the playing game experiments
Figure 4: Playing minigames.
For this task, we evaluated GPT-3.5 and GPT-4. For each model, we tried two settings: the first is to condition the LLM on the walkthrough like LABEL:lst:walkthrough; the second is to include in the prompt the information about the nearby locations, and an example of the full prompt is given in LABEL:lst:playgameprompt. The information about nearby locations is the ground-truth information that the LLM, in principle, should have learned from the walkthrough prefix. If the LLM had a perfect mapping and navigation ability, it would be able to perfectly spell it out and use that information to guide its decision making. Figure 4 presents the results of this experiment. GPT-4 significantly outperforms GPT-3.5 in playing these minigames, consistent with their relative performance when answering the DF and RF questions of our MANGO benchmark. For each of the GPT models, having access to nearby location information significantly improves its performance, demonstrating that a strong mapping and navigation ability is essential to succeeding at relevant downstream tasks.
4 Related Work
Over the past few years, the field of natural language processing has experienced remarkable advancements with the emergence of large language models. This progress has spurred a multitude of research endeavors that propose benchmarks challenging the limits of these models. Those benchmarks assess the capacities of LLMs in linguistics (Wang et al., 2018; 2019), reading comprehension (Richardson et al., 2013; Lai et al., 2017), commonsense reasoning (Zellers et al., 2019; Bisk et al., 2020; Huang et al., 2019; Talmor et al., 2019), arithmetic reasoning (Miao et al., 2020; Cobbe et al., 2021; Patel et al., 2021), and knowledge memorization and understanding (Clark et al., 2018; Mihaylov et al., 2018; Khot et al., 2020; Clark et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2022). Recent models have achieved remarkable performance not only on these benchmarks, but also across a diversity of human-oriented academic and professional exams (OpenAI, 2023) as well as general tasks (Bubeck et al., 2023). Our benchmark presents a unique challenge to large language models, evaluating their capacity to acquire spatial knowledge about new environments and answering complex navigation questions; it is a dimension orthogonal to the aforementioned reasoning abilities.
The advances of LLMs have sparked a recent wave of endeavors that integrate these models into embodied agents (Huang et al., 2022c; Yang et al., 2023; Vemprala et al., 2023; Wang et al., 2023a). Generally, they utilize language models as a means to understand human instructions and plan executable actions (Driess et al., 2023; Liang et al., 2022; Huang et al., 2022b; Ichter et al., 2023). This includes instructions related to object manipulation and tool operation (Wang et al., 2023b; Ren et al., 2023) as well as localization and navigation (Majumdar et al., 2020; Gadre et al., 2023; Shah et al., 2023; Huang et al., 2022a). Our MANGO benchmark aligns with the growing trend to deploy LLMs in embodied agents and provides a comprehensive investigation of their capacities in mapping and navigation. Our benchmark operates in text-based environments, distinguishing itself from previous benchmarks (Puig et al., 2018; Shridhar et al., 2020; Fan et al., 2022) that allow agents to utilize visual signals. This “text-only” design enables us to conduct controlled experiments that investigate the capacity of language models to acquire knowledge about environments solely from textual inputs and answer navigation questions based on that knowledge. It complements the existing benchmark and methodological research in vision-language navigation (Duvallet et al., 2014; Mei et al., 2016; Anderson et al., 2017; Fried et al., 2018; Zhu et al., 2020; Min et al., 2021). Our work is related to recent studies that demonstrate the emergence of maps with learned neural representations as a consequence of navigation (Huynh et al., 2020; Wijmans et al., 2023) with the key distinction that our agents are provided with textual descriptions of their environments.
Given our focus on mapping and navigation, it is worth noting the work on simultaneous localization and mapping (SLAM), a classic problem in which a mobile agent (e.g., a robot or hand-held camera) is tasked with mapping an a priori unknown environment while concurrently using its estimated map to localize itself in the environment (Mur-Artal et al., 2015; Cadena et al., 2016). Particularly relevant are the methods that maintain spatial-semantic maps of the environments based on natural language descriptions (Walter et al., 2013; Hemachandra & Walter, 2015), however they rely on non-linguistic observations (e.g., vision) to ground these descriptions.
5 Conclusion
We present MANGO, a benchmark that evaluates the mapping and navigation abilities of large language models. Our benchmark covers a diversity of mazes as well as a variety of evaluation and analysis programs, offering a comprehensive testbed in a great breadth and depth. In our experiments, the current best model still performs poorly on the benchmark, with a sharp degradation on the more difficult questions. We release our benchmark—along with the source code for data generation and evaluation—to track the advances of the mapping and navigation capabilities of future LLMs as well as to facilitate future research in related areas. Several interesting future directions are discussed in Appendix D.
Acknowledgments
This work was supported by a research gift to the last author by Adobe Research. It was also supported by the TRI University 2.0 Project, the U.S. National Science Foundation (Grant Number CNS-1948457 and CNS-2340171), and Office of Advanced Scientific Computing Research at Department of Energy (Award Number DE-SC0024424). The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. We thank the anonymous COLM reviewers and meta-reviewer for their constructive feedback. We also thank our colleagues at TTIC and UChicago for helpful discussion.
Appendix A Benchmark Details
Our data and program is released at https://mango.ttic.edu and https://github.com/oaklight/mango/. In the data folder, each game has a folder that contains multiple JSON files. The most important files are the DF and RF skeletons (where X is the game name such as zork or detective):
  • X.df: it contains the DF skeletons like LABEL:lst:dfqraw.
  • X.rf: it contains the RF skeletons like LABEL:lst:rfqraw.
As introduced in section 2.2, each skeleton is paired with an ANSWERABLE label and an EASY label: given a prefix of the walkthrough, the ANSWERABLE label indicates whether this skeleton is answerable, and the EASY label will decide whether this skeleton is easy or hard given this walkthrough prefix. Table 3 displays some important statistics of the full dataset broken down into each individual maze, including the number of DF and RF skeletons. In total, our full dataset has about 3M DF questions and 200K RF questions. On average, each maze has around 60K DF questions and 4K RF questions. Noticeably, the 53 games in our dataset are very diverse: they cover a range of topics and genres; they are situated in different eras; they cover a wide variety of maps (small vs. big houses, long vs. short halls, towns vs. forests, verbose vs. concise scene descriptions, etc). In addition, we also provide the following data files for easy reference:
  • X.walkthrough: it contains the full walkthrough of the game. See section A.1 for details.
  • X.locations: it lists all the locations. Details about annotating these locations can be found in section A.2.
  • X.moves: it lists all the moves that may change the location.
  • X.all_pairs: it contains all the pairs of distinct locations.
  • X.all2all: it contains all the simple paths between any pair of distinct locations.
Maps # Locs # Edges Avg Len Path # Steps DF RF
Easy Hard Easy Hard
905 5 7 1.88 22 11 5 11 5
advent 70 137 19.38 277 24788 4220 4694 0
adventureland 23 48 7.46 170 1117 208 484 22
afflicted 13 24 3.18 99 156 0 156 0
anchor 118 228 27.46 531 66133 23564 13573 0
awaken 15 28 5.02 57 365 45 171 25
balances 12 20 3.22 122 136 1 94 1
ballyhoo 39 90 9.26 416 18148 6298 1444 0
curses 150 304 33.83 816 1951931 549433 22201 0
cutthroat 52 108 10.76 336 5969 2688 1832 59
deephome 66 127 8.64 327 4161 235 4161 1
detective 32 41 8.76 51 505 6 505 6
dragon 24 52 8.9 101 1100 2262 529 23
enchanter 57 110 10.93 265 5118 2602 2773 99
enter 18 34 3.88 102 306 0 306 0
gold 22 44 4.99 345 682 0 462 0
hhgg 42 65 7.02 361 3794 3 1602 1
hollywood 109 219 27.33 397 21641 15557 11451 213
huntdark 12 13 4.24 67 66 2 66 2
infidel 58 122 9.27 250 2786 1729 2158 453
inhumane 43 91 6.49 122 1623 3020 1474 332
jewel 43 74 8.66 223 1157 61 1157 61
karn 56 124 15.36 362 24479 108983 3025 0
library 7 12 2.48 52 42 0 42 0
loose 12 21 4.18 50 94 27 94 27
lostpig 13 26 4.5 146 492 105 114 0
ludicorp 86 176 14.96 364 7099 7281 6002 1308
lurking 60 116 13.04 294 5187 2316 2946 105
moonlit 6 9 2.2 59 18 7 18 7
murdac 84 157 11.05 304 6914 2016 5967 1005
night 20 41 6.93 90 633 59 380 0
omniquest 32 65 7.55 78 648 1431 642 90
partyfoul 4 8 1.92 56 24 0 12 0
pentari 18 31 3.76 49 208 20 208 4
planetfall 69 138 12.41 399 7100 2246 3887 247
plundered 45 87 13.46 189 3705 2394 1393 26
reverb 17 31 5.26 74 321 20 256 16
seastalker 20 43 7.55 204 1123 1143 260 3
sherlock 71 140 12.39 339 16417 11758 4214 27
snacktime 4 6 1.5 34 12 0 12 0
sorcerer 64 120 11.72 254 4042 2285 2503 264
spellbrkr 73 111 13.62 412 8713 1762 3462 1
spirit 229 466 26.39 1264 111102 47410 51531 226
temple 24 46 6.07 181 563 106 486 21
theatre 20 38 5.33 296 329 130 329 51
trinity 94 189 20.39 610 65049 131324 8742 0
tryst205 73 139 10.04 518 5089 392 3983 181
wishbringer 46 94 13.32 184 4988 5646 2070 0
yomomma 9 25 3.55 98 189 263 64 0
zenon 17 32 4.85 83 144 128 144 128
zork1 84 166 20.79 396 20984 26857 6889 0
zork2 66 136 11.79 296 15174 13081 3798 178
zork3 55 105 15.37 273 6202 2754 2294 133
ztuu 16 30 3.27 84 225 15 225 15
Table 3: Statistics of full data. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks. When counting the easy and hard skeletons, we assume that the full walkthrough will be used. For the statistics of the data used in our experiments, please see Tables 5, 8, 7 and 6.
In the following subsections, we will explain the details about collecting this data.
A.1 Walkthrough Details
In this section, we document the technical details about walkthroughs.
The only game that doesn’t have a walkthrough is Leather Goddesses of Phobos (LGoP).
There may be multiple correct ways to complete a game with the same level of efficiency. So the program has some randomness. In our experiments, we fixed the random seed for better reproducibility.
Each action in the program-generated walkthroughs is a highly abbreviated symbol such as E for East and NW for Northwest. For a better readability, we use the full words in our enhanced walkthrough such as LABEL:lst:walkthrough.
A.2 Location Resolution Details
For each maze, our human annotator read the walkthrough and annotated all the locations. In most cases, the surrounding description given by the game engine includes the name of the location, and the annotator needs to manually extract it from the text; this process is difficult to automate because the text is often unstructured and an automatic extractor is hard to build. What makes this annotation tricky is
  • a location may be visited more than once in the walkthrough but we should avoid assigning multiple names to it.
  • distinct locations may be referred to in the same way by the game engine but we should distinguish them. The game of Night is an example: hallways on different floors are all referred to as Hall; we renamed each of them, e.g., with Hall (1st floor, north end).
We solved the problems above by hacking into the source code of the game engines and conducting multiple rounds of human verification. First, for each maze, we checked the source code of the game provided by the Jericho game suite (Hausknecht et al., 2020) and found the unique ID for each location. Matching IDs with human annotations allows us to perform the following post processing:
  • when a location is given multiple names, our human annotators work together to select the most proper unique name that they all agree on. The selection principles are: it is descriptive; whenever this location is visited in the walkthrough, the name has an intuitive match with the surrounding description given by the game engine.
  • when multiple locations share a name, our human annotators work together to distinguish them by adding descriptive marks. An example is the Halls mentioned above: we renamed them to be Hall (1st floor, north end of north/south hall), Hall (1st floor, middle of north/south hall), and Hall (2nd floor, middle of north/south hall). There are rare cases in which all the annotators agreed that no marks could be added and the location names had to be kept fuzzy (i.e., a name corresponds to multiple different locations). The rationale is: if a human may confuse with these locations, then it is reasonable for a model to have the same confusion. Then allowing them to share the name is essentially to apply a looser evaluation to the models: e.g., if the name Forest is overloaded, then when the model answers “how to reach Forest from House”, any path that ends at any of the Forests will be considered to be correct. This treatment is equivalent to merging the locations with the same human-annotated name.
In our repository, there is a data-intermediate folder that tracks such intermediate annotations. In the folder of each game, the JSON files anno2code and code2anno track the mapping between machine IDs and human-annotated location names.
Why don’t we just use the unique IDs as the location names? Because the IDs are often not intuitive or descriptive and they are often just strings of digits. Such IDs may not match any content in the walkthrough so a human or model may be confused when asked about the path between “loc12” and “loc5” after reading the very descriptive walkthrough.
In addition, we also provide an alternative version of our data in which location names include machine IDs. Precisely, while resolving each location name, we
  • add the machine ID (e.g., “(obj59)”) to the location name as a mark (e.g., “in debris room” to “in debris room (obj59)”);
  • if the location name already has a mark (e.g., 1st floor, north end of north/south hall, thanks to the name resolution process), we replace the mark with the machine ID (e.g., “hall (1st floor, middle of north/south hall)” to “hall (obj66)”).
A.3 Move Resolution Details
During human annotation, we labeled all the moves that change the locations. Like in section A.2, we used the source code of the game engine to verify the human annotations. This annotation led to a map for each game—like what’s shown in Figure 1—but this map is incomplete since there exist implicit moves between locations. For example, south is a movement that can end up at Altar if we start from Temple; see Figure 1. It means that north is also a possible move from Altar and it leads to Temple; but this edge has never explicitly shown up in the walkthrough. We would like our questions to cover such implicit edges, so we examined every possible implicit edge and inserted the valid ones into our map (though Figure 1 only displays the explicit edges).
The examination was carried out through real game playing by our human annotators. For each directional move (e.g., south), we tested if its reverse directional move (e.g., north for south, down for up) could lead to the previous location. Not every move is directional: e.g., in Zork-I, you may enter and exit the House; pray moves the player from Altar to Forest. We had such edges examined by human annotators: for some (e.g., enter), we could find intuitive reverse moves and verify them; for the others (e.g., pray), we didn’t propose any reverse. Usually, the list of moves we ended up for a game includes eight possible directional moves as well as a few special moves.
A.4 Path Details
Once we have figured out all the locations and moves for a game, we will end up with a map. We collected all the unique pairs of distinct locations in the map and stored them in the all_pairs file: for each pair of locations A and B, we could ask a route finding question that aims to reach B from A. Note that A,B and B,A are different pairs.
For each pair of the starting point A and destination B, we collected all the simple paths P that connects from A to B. Each (A,B,P) tuple defines a destination finding question about where one will be if they go through path P from A.
A.5 Program Details
Our graph and path operations (e.g., finding simple paths) are handled by the networkx package. Its documentation is at https://networkx.org/. Particularly, the program that finds all the simple paths for a pair of graph nodes is at https://networkx.org/documentation/stable/reference/algorithms/generated/networkx.algorithms.simple_paths.all_simple_paths.html.
A.6 Evaluation Details
Another important—yet more strict—evaluation is the reasoning accuracy. This evaluation requires the language model to spell out its planned trajectories of moves when answering questions (like requested in our LABEL:lst:maptemplate and LABEL:lst:navtemplate). For an RF question, the reasoning process is correct if and only if the model-generated trajectory is a valid path from the starting position S to the destination D: the first step starts from S; each step starts from where it ended up in the last step; the final step reaches D. For a DF question, the model-generated trajectory has to be a valid path from S to the model-generated destination D; in addition, the sequence of moves in the trajectory has to match the given list of actions A. Like we explained in section 2.3, the “match” here is not an exact match: if the closest valid move is the correct move, then it is counted as a “match”.
Appendix B Experiment Details
In this section, we present our experiment details for reproducing the results.
B.1 Model Configuration Details
The specific GPTs used in our experiments are GPT-4-0314 and for GPT-3.5-turbo-0301. For Claude-1 and Claude-2, the specific model versions we are using are Claude-instant-1.2 and Claude-2.0. For 14B RWKV model, the specific model version we are using is “RWKV-4-Pile-14B-20230313-ctx8192-test1050”. For Llama-2, we used the ckeckpoint officially released by Meta.
For all the models, we set the temperature of the LLMs to be 0 for reproducibility.
B.2 Prompt Details
For GPTs, our prompts are the concatenation of walkthrough like in LABEL:lst:walkthrough and the questions like in LABEL:lst:dfq and LABEL:lst:rfq; the questions are obtained by filling the templates in LABEL:lst:maptemplate and LABEL:lst:navtemplate. When calling OpenAI API, we set the “role” to be “user” and the “content” to be the prompt. After receiving the response from the API call, we processed the output string by fetching the content from the structured output (recall that we request the models to return Python lists of Python dictionaries).
What prompts are to LLMs are like what hyperparameters are to classifical deep neural nets. Our experiments show that the structure of LLM output is sensitive to the prompts, and thus we had to carefully tune the prompts such that LLMs could return easy-to-parse output. For example, we found it helpful to ask LLMs to format their answers as a Python list of Python dictionaries; we also found it helpful to end our prompt with the symbol ‘[’—i.e., the start symbol of a Python list—to elicit the LLM to actually output in the desired format. Fortunately, as long as the output follows the desired structure, its content is relatively robust to the prompts. In our pilot experiments, we found that different prompts (created by different authors) would generate the same or similar content (regardless of its format) and achieve the same level of success rates (after parsing).
To ensure that we could obtain the optimal results of RWKV and Llamas, we tuned the prompts—more precisely, experimented with variants of the prompts of GPTs—and ended up with a set of new prompts that are mostly the same but exhibit some prose differences. For example, at the beginning of the prompts, we added “Here is a walkthrough of a text game.” It is worth noting that, even though we carefully tuned the prompts, the non-GPT models still suffer a high chance of failing to return well-structured answers. Table 4 shows the average number of answered questions as well as how many of them received ill-formatted answers. As we could see, GPT-3.5 and GPT-4 could generate well-structured answers for a large portion of the DF and RF questions, but the other models often gave ill-structured answers. As a result, we could only evaluate the non-GPT models on the questions to which the answers were well-structured and thus could be parsed. We also experimented with the function-calling interface of GPTs, but it didn’t lead to an increased amount of well-structured answers compared to our prompt design.
Model RF Question DF Question
# answerable # ill-structured # answerable # ill-structured
RWKV 277.32 240.04 397.08 325.98
Llama-2 138.40 112.25 157.98 98.0
Claude-1 184.34 3.88 244.92 4.66
Claude-2 184.34 6.28 244.92 1.86
GPT-3.5 184.34 30.81 244.92 65.60
GPT-4 184.34 2.49 244.92 5.74
Table 4: Average (per-maze) numbers of answerable questions and ill-structured answers for each model.
B.3 Walkthrough and Question Details
Due to the context window size of an LLM, it is often the case that we have to only use a prefix of the walkthrough when evaluating an LLM on a maze. Recall from section 2.2 that not every question is answerable given a walkthrough prefix. Therefore, each model was evaluated on a different set of questions. Tables 5, 8, 7 and 6 display the statistics about the data that each LLM was actually evaluated on in our experiments.
Maps # Locs # Edges Avg Len Path # Steps DF RF
Easy Hard Easy Hard
905 5 7 1.88 21 11 5 11 5
advent 31 57 7.79 70 692 100 532 100
adventureland 18 35 6.13 70 579 80 260 46
afflicted 10 16 2.78 40 67 0 67 0
anchor 13 24 3.94 24 95 91 82 74
awaken 14 24 4.8 44 262 12 157 12
balances 11 18 3.09 67 96 8 76 8
ballyhoo 14 28 4.8 56 225 99 156 13
curses 14 24 3.3 53 122 13 122 0
cutthroat 22 40 5.81 62 303 158 202 107
deephome 17 28 4.35 49 175 10 175 10
detective 26 34 7.17 43 334 4 334 4
dragon 14 25 3.59 29 105 64 111 58
enchanter 21 38 5.69 53 216 165 216 165
enter 2 1 1.0 20 1 0 1 0
gold 11 17 2.82 47 83 0 83 0
hhgg 8 9 2.6 51 29 1 29 1
hollywood 8 14 2.71 50 43 13 43 13
huntdark 10 9 3.67 55 45 0 45 0
infidel 13 26 3.6 55 114 138 88 68
inhumane 21 40 4.8 49 275 240 261 159
jewel 16 30 4.39 60 166 74 166 74
karn 19 35 6.37 65 339 86 231 63
library 7 12 2.48 49 42 0 42 0
loose 8 14 3.0 39 56 0 56 0
lostpig 6 9 1.96 56 16 9 16 9
ludicorp 22 43 4.91 70 351 111 351 111
lurking 10 16 2.89 56 66 16 65 16
moonlit 4 6 1.67 45 9 3 9 3
murdac 30 52 6.34 70 537 195 528 183
night 20 41 6.93 70 633 59 380 0
omniquest 29 59 7.75 70 536 1198 290 298
partyfoul 4 6 1.67 24 12 0 11 1
pentari 18 30 3.72 48 208 4 208 4
planetfall 21 37 5.27 68 246 50 246 50
plundered 10 11 3.23 32 47 0 47 0
reverb 12 21 4.07 40 129 12 120 12
seastalker 10 15 2.7 53 50 3 50 3
sherlock 8 13 2.48 33 43 5 43 0
snacktime 4 6 1.5 33 12 0 12 0
sorcerer 15 26 3.53 54 120 17 120 17
spellbrkr 15 23 3.6 52 120 17 120 16
spirit 15 26 3.55 50 171 12 171 12
temple 10 15 3.11 46 48 18 48 18
trinity 10 10 3.61 45 45 1 45 1
tryst205 9 14 1.93 65 57 0 57 0
wishbringer 18 30 5.2 45 174 48 175 47
yomomma 8 15 2.39 41 49 17 31 18
zenon 13 24 4.05 63 83 73 83 73
zork1 19 34 7.16 70 351 46 279 45
zork2 22 33 5.65 50 242 18 146 103
zork3 23 44 6.84 61 600 167 394 68
ztuu 11 18 2.79 43 91 0 91 0
Table 5: Map statistics for GPTs and Claudes. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks.
Maps # Locs # Edges Avg Len Path # Steps DF RF
Easy Hard Easy Hard
905 5 7 1.88 21 11 5 11 5
advent 31 57 7.79 70 692 100 532 100
adventureland 18 35 6.13 70 579 80 260 46
afflicted 11 20 2.95 70 100 10 100 10
anchor 25 46 5.99 70 327 153 302 132
awaken 15 28 5.02 56 365 45 171 25
balances 11 18 3.09 70 96 8 76 8
ballyhoo 17 35 4.83 70 302 188 213 59
curses 14 27 3.62 70 182 13 182 0
cutthroat 25 49 6.66 70 471 362 360 216
deephome 27 49 4.83 70 429 19 429 19
detective 32 40 8.79 50 505 4 505 4
dragon 21 44 7.11 70 533 990 272 148
enchanter 23 43 6.35 70 265 219 265 219
enter 14 26 3.36 70 117 65 117 65
gold 15 25 3.45 70 143 0 143 0
hhgg 9 11 2.85 70 38 1 38 1
hollywood 12 22 3.36 70 84 48 84 48
huntdark 12 11 4.33 66 66 0 66 0
infidel 24 48 7.53 70 312 446 264 288
inhumane 30 57 5.54 70 614 555 483 280
jewel 17 32 4.4 70 187 85 187 85
karn 19 35 6.37 70 339 86 231 63
library 7 12 2.48 51 42 0 42 0
loose 12 21 4.18 49 94 27 94 27
lostpig 7 11 2.28 70 22 14 22 14
ludicorp 22 43 4.91 70 351 111 351 111
lurking 16 29 4.29 70 144 97 143 97
moonlit 6 9 2.2 58 18 7 18 7
murdac 30 52 6.34 70 537 195 528 183
night 20 41 6.93 70 633 59 380 0
omniquest 29 59 7.75 70 536 1198 290 298
partyfoul 4 9 1.97 55 24 6 11 1
pentari 18 30 3.72 48 208 4 208 4
planetfall 22 39 5.48 70 267 63 267 63
plundered 22 37 6.02 70 450 52 289 26
reverb 17 31 5.26 70 321 20 253 19
seastalker 10 15 2.7 70 50 3 50 3
sherlock 18 28 4.36 70 175 5 175 0
snacktime 4 6 1.5 33 12 0 12 0
sorcerer 26 46 7.09 70 340 48 340 48
spellbrkr 20 31 4.84 70 295 23 276 21
spirit 22 41 4.09 70 354 87 354 87
temple 19 33 4.72 70 178 69 178 69
trinity 17 17 5.96 70 136 1 136 1
tryst205 9 15 1.94 70 64 0 64 0
wishbringer 21 40 6.34 70 259 214 251 169
yomomma 9 20 2.74 70 82 59 43 21
zenon 14 26 4.27 70 96 86 96 86
zork1 19 34 7.16 70 351 46 279 45
zork2 22 45 7.01 70 536 754 239 130
zork3 23 45 6.93 70 627 174 414 70
ztuu 15 26 3.15 70 183 0 183 0
Table 6: Map statistics for RWKV, RWKV-S, Llama-1-S, Llama-2-S, Code-Llama-S, Code-Llama-Instruct-S. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks.
Maps # Locs # Edges Avg Len Path # Steps DF RF
Easy Hard Easy Hard
905 5 7 1.88 19 11 5 11 5
advent 12 21 3.73 18 67 30 67 30
adventureland 8 11 2.09 20 32 1 32 1
afflicted 3 3 1.25 8 4 0 4 0
anchor 5 5 1.91 5 10 1 7 4
awaken 6 8 2.32 10 24 4 21 4
balances 5 5 1.91 16 10 1 10 1
ballyhoo 5 4 2.0 9 10 0 10 0
curses 6 8 1.95 12 17 2 17 0
cutthroat 2 2 1.0 13 1 1 1 1
deephome 3 3 1.25 11 3 1 3 1
detective 7 8 2.35 16 23 0 23 0
dragon 4 4 1.57 4 6 1 6 1
enchanter 9 13 2.63 15 38 13 38 13
enter 0 0 0 3 0 0 0 0
gold 5 5 1.91 9 11 0 11 0
hhgg 4 4 1.57 15 6 1 6 1
hollywood 4 5 1.56 10 6 3 6 3
huntdark 3 2 1.33 10 3 0 3 0
infidel 6 8 1.9 9 16 5 17 4
inhumane 6 8 1.95 14 19 2 19 2
jewel 6 9 2.08 17 21 4 21 4
karn 3 4 1.33 20 6 0 6 0
library 2 1 1.0 12 1 0 1 0
loose 4 3 1.67 8 6 0 6 0
lostpig 3 3 1.25 17 3 1 3 1
ludicorp 8 11 2.62 24 32 0 32 0
lurking 3 3 1.25 21 4 0 4 0
moonlit 3 2 1.33 14 3 0 3 0
murdac 11 19 3.75 25 110 8 102 8
night 11 12 3.62 17 58 0 58 0
omniquest 10 18 2.73 22 54 36 54 36
partyfoul 0 0 0 4 0 0 0 0
pentari 8 10 2.78 7 28 4 28 4
planetfall 2 2 1.0 20 1 1 1 1
plundered 2 1 1.0 9 1 0 1 0
reverb 5 5 1.91 8 10 1 10 1
seastalker 2 2 1.0 12 1 1 1 1
sherlock 4 3 1.67 7 6 0 6 0
snacktime 2 2 1.0 15 2 0 2 0
sorcerer 4 7 1.56 16 7 2 7 2
spellbrkr 3 3 1.25 10 3 1 3 1
spirit 4 4 1.57 9 6 1 6 1
temple 3 3 1.25 11 3 1 3 1
trinity 4 4 1.57 14 6 1 6 1
tryst205 3 2 1.33 9 3 0 3 0
wishbringer 7 8 2.35 11 23 0 23 0
yomomma 4 5 1.57 8 6 1 4 3
zenon 4 4 1.57 21 6 1 6 1
zork1 8 12 2.27 22 34 3 34 3
zork2 8 10 2.65 11 29 2 29 2
zork3 7 14 2.59 17 40 24 30 12
ztuu 5 4 2.0 5 10 0 10 0
Table 7: Map statistics for Llama-1. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks.
Maps # Locs # Edges Avg Len Path # Steps DF RF
Easy Hard Easy Hard
905 5 7 1.88 21 11 5 11 5
advent 25 45 7.52 54 473 76 355 76
adventureland 18 33 5.91 56 413 42 260 46
afflicted 9 14 2.43 32 46 0 46 0
anchor 9 15 3.0 19 50 16 41 13
awaken 13 18 3.98 34 114 4 90 4
balances 8 12 2.51 49 38 1 38 1
ballyhoo 12 21 3.98 43 111 35 101 11
curses 10 17 2.93 42 67 9 67 0
cutthroat 19 28 4.75 42 180 9 136 53
deephome 12 21 3.43 40 102 10 102 10
detective 25 32 6.99 40 307 4 307 4
dragon 11 20 3.16 22 68 42 70 40
enchanter 17 31 4.61 41 142 114 142 114
enter 0 0 0 15 0 0 0 0
gold 8 14 2.5 40 56 0 56 0
hhgg 7 8 2.35 39 22 1 22 1
hollywood 7 8 2.5 34 21 3 21 3
huntdark 7 6 2.67 41 21 0 21 0
infidel 12 24 3.34 42 102 116 78 54
inhumane 18 34 4.82 40 218 168 190 116
jewel 15 25 3.77 46 141 30 141 30
karn 15 27 4.11 49 136 38 144 30
library 7 11 2.22 40 32 0 32 0
loose 8 13 2.86 28 49 0 49 0
lostpig 5 7 1.69 47 11 5 11 5
ludicorp 19 37 4.14 65 291 51 291 51
lurking 6 9 2.42 40 30 1 29 1
moonlit 3 2 1.33 36 3 0 3 0
murdac 23 43 5.9 62 355 171 346 160
night 20 39 5.79 53 380 19 380 0
omniquest 26 49 6.28 58 348 180 222 192
partyfoul 4 6 1.67 19 12 0 11 1
pentari 17 29 3.67 42 191 4 191 4
planetfall 17 29 4.96 52 164 16 164 16
plundered 7 7 2.59 26 22 0 22 0
reverb 12 15 3.54 28 68 2 68 2
seastalker 10 15 2.7 43 50 3 50 3
sherlock 6 9 2.08 21 25 0 25 0
snacktime 4 6 1.5 33 12 0 12 0
sorcerer 10 18 2.3 44 60 13 60 13
spellbrkr 10 13 2.71 41 60 2 59 2
spirit 13 21 3.22 38 99 6 99 6
temple 9 14 2.84 37 39 18 39 18
trinity 8 8 2.93 37 28 1 28 1
tryst205 8 13 1.92 49 49 0 49 0
wishbringer 18 30 5.2 38 174 48 175 47
yomomma 7 13 2.18 31 36 14 26 10
zenon 12 22 3.71 52 70 62 70 62
zork1 19 32 7.25 56 332 22 279 45
zork2 19 27 5.15 42 176 18 107 76
zork3 18 35 4.78 48 282 100 199 45
ztuu 8 11 2.26 28 38 0 38 0
Table 8: Map statistics for Llama-2, Code-Llama, Code-Llama-Instruct. Here, # Locs represents the number of locations in each walkthrough; # Edges represents the number of edges; AVG LEN PATH denotes the average length of all paths; # STEPS indicates the number of steps in each walkthrough; EASY and HARD of the DF and RF respectively represent the number of easy and hard skeletons of the DF and RF tasks.
Appendix C More Results
C.1 Success Rate Weighted by Route Length
Following the evaluation standard proposed by Anderson et al. (2018), we also computed the success rates weighted by route length, which are shown in Figure 5. As we can see, the ranking of the models stays the same as in Figure 2.
(a) On easy (left) and hard (right) DF questions.
(b) On easy (left) and hard (right) RF questions.
Figure 5: Success rates weighted by route length on DF (5(a)) and RF (5(b)) questions, averaged over all $53$ mazes.
C.2 Reasoning Accuracy Results
In this section, we present the results of each LLM measured by reasoning accuracy, the metric introduced in section A.6. Results are shown in Figure 6, with their pair-wise comparison shown in Table 9. As we can see, the trend measured by this metric is similar to what’s shown in section 3: GPT-4 is the best among all the evaluated models but still suffers a low accuracy. Figure 7 shows the reasoning accuracies of GPT-3.5 vs. GPT-4 broken down into individual games, showing similar patterns with Figure 3.
(a) On easy (left) and hard (right) DF questions.
(b) On easy (left) and hard (right) RF questions.
Figure 6: Reasoning accuracy of each model on DF (6(a)) and RF (6(b)) questions, averaged over all $53$ mazes.
Method RWKV Llama-2 Claude-1 Claude-2 GPT-3.5 GPT-4 $\overline{{\rm HARD}}|$
RWKV * 0.00 | 0.02 0.00 | 0.15 0.00 | 0.18 0.00 | 0.05 0.00 | 0.41 *
Llama-2 0.15 | 0.00 * 0.01 | 0.20 0.01 | 0.24 0.01 | 0.07 0.01 | 0.52 *
Claude-1 0.51 | 0.00 0.62 | 0.15 * 0.11 | 0.13 0.13 | 0.05 0.11 | 0.29 *
Claude-2 0.60 | 0.00 0.72 | 0.15 0.58 | 0.47 * 0.16 | 0.05 0.13 | 0.29 *
GPT-3.5 0.32 | 0.00 0.38 | 0.15 0.28 | 0.52 0.28 | 0.63 * 0.05 | 0.35 *
GPT-4 0.71 | 0.00 0.79 | 0.15 0.67 | 0.47 0.67 | 0.58 0.73 | 0.28 * *
|$\underline{{\rm EASY}}$
(a) Pairwise comparison on easy (lower left) and hard (higher right) DF questions.
Method RWKV Llama-2 Claude-1 Claude-2 GPT-3.5 GPT-4 $\overline{{\rm HARD}}|$
RWKV * 0.00 | 0.00 0.00 | 0.12 0.00 | 0.20 0.00 | 0.03 0.00 | 0.51 *
Llama-2 0.02 | 0.00 * 0.00 | 0.14 0.00 | 0.19 0.00 | 0.04 0.00 | 0.45 *
Claude-1 0.33 | 0.00 0.32 | 0.02 * 0.10 | 0.17 0.12 | 0.02 0.10 | 0.42 *
Claude-2 0.46 | 0.00 0.45 | 0.03 0.45 | 0.31 * 0.18 | 0.02 0.17 | 0.43 *
GPT-3.5 0.15 | 0.00 0.16 | 0.03 0.15 | 0.34 0.14 | 0.48 * 0.02 | 0.46 *
GPT-4 0.56 | 0.00 0.55 | 0.02 0.53 | 0.31 0.53 | 0.45 0.56 | 0.15 * *
|$\underline{{\rm EASY}}$
(b) Pairwise comparison on easy (lower left) and hard (higher right) RF questions.
Table 9: Reasoning accuracies on DF and RF questions broken down into pairwise comparison.
(a) GPT-3.5 vs. GPT-4 on DF questions.
(b) GPT-3.5 vs. GPT-4 on RF questions.
Figure 7: Reasoning accuracies of GPT-3.5 and GPT-4 broken down into individual games. Similar to Figure 3, in each subfigure, the left scatterplot is for easy questions while the right is for hard questions.
C.3 Analysis Details of GPTs
As discussed in section 3.3, we used regression analysis to understand the effects of the aforementioned variables on model performance. In particular, for each model (GPT-3.5 or GPT-4) on each type of question (DF or RF, easy or hard), we ran single-variable linear regression to understand how the success rate varies with each of the variables of interest. Table 10 displays the regression results.
Metric GPT-3.5 GPT-4
DF RF DF RF
$\beta$ $p$ $\beta$ $p$ $\beta$ $p$ $\beta$ $p$
# locations $-0.076$ $0.001$ $-0.084$ $0.000$ $-0.073$ $0.000$ $-0.120$ $0.000$
# exp edges $-0.068$ $0.003$ $-0.081$ $0.001$ $-0.067$ $0.000$ $-0.096$ $0.003$
# conf locations $-0.056$ $0.017$ $-0.067$ $0.006$ $-0.065$ $0.000$ $-0.088$ $0.006$
avg len easy $-0.066$ $0.005$ $-0.081$ $0.001$ $-0.066$ $0.000$ $-0.131$ $0.000$
avg len scene $-0.037$ $0.119$ $-0.002$ $0.939$ $0.031$ $0.036$ $0.069$ $0.035$
(a) Regression analysis results on easy questions.
Metric GPT-3.5 GPT-4
DF RF DF RF
$\beta$ $p$ $\beta$ $p$ $\beta$ $p$ $\beta$ $p$
# locations $-0.049$ $0.088$ $-0.055$ $0.043$ $-0.083$ $0.007$ $-0.115$ $0.006$
# exp edges $-0.052$ $0.070$ $-0.064$ $0.018$ $-0.080$ $0.010$ $-0.107$ $0.011$
# imp edges $-0.059$ $0.038$ $-0.032$ $0.252$ $-0.118$ $0.000$ $-0.152$ $0.000$
# conf locations $-0.045$ $0.125$ $-0.055$ $0.043$ $-0.068$ $0.030$ $-0.081$ $0.061$
avg len hard $-0.072$ $0.011$ $-0.055$ $0.046$ $-0.094$ $0.002$ $-0.122$ $0.004$
avg # imp in hard $-0.057$ $0.046$ $-0.022$ $0.428$ $-0.081$ $0.009$ $-0.096$ $0.024$
avg len scene $0.032$ $0.268$ $0.050$ $0.069$ $0.084$ $0.007$ $0.105$ $0.013$
(b) Regression analysis results on hard questions.
Table 10: Regression analysis results, where $\beta$ is the regression coefficient and $p$ denotes the $p$-value. When $p<0.001$, we write $0.000$ for presentation simplicity.
In our pilot experiments, we ran a multivariate regression analysis that used the aforementioned variables jointly. However, the results of this regression are misleading: due to colinearity among the explanatory variables, the estimated coefficients are unreliable and the $p$-values are inflated. We also tried principled component regression, but the first principle component has nearly equal loadings across all the variables, making it inconvenient to interpret the results.
C.4 Results of Search-Based Approaches
We also explored an approach that first maps a walkthrough to a symbolic graph and then uses a search algorithm to answer the given (DF or RF) question. However, we found it to be very challenging to translate natural language walkthrough into searchable symbolic graphs in the first place, making this approach not promising.
We tried two different settings. In the first setting, we did 5-shot prompting and each in-context example is a single S-A-D triplet; the examples are the first five steps in the walkthrough. In the second setting, we did 0-shot prompting and the prompt includes the entire walkthrough; the LLM has to generate a sequence of S-A-D triplets which then could be used to construct the graph. For each setting, we ran experiments on the following six mazes: Zork-I, Night, Partyfoul, Plundered, Spirit, and Temple.
In both settings, GPT-3.5 yields a low accuracy of the S-A-D triplet completion, indicating the difficulty of translating natural language into symbolic graphs. In the first setting, the average accuracy is 70.5 %. In the second setting, the average success rate is 6.3 %. The second setting is more challenging because any mistake would cause all subsequent steps to be incorrect. Some LLM mistakes and our error analysis can be found at https://github.com/Oaklight/mango/tree/camera-ready/utils/supp_exp.
C.5 More Llama Results
The Llama-2 we used in the experiments of section 3 is the base model. We also experimented with the the 32.5B Llama-1 model released earlier (Touvron et al., 2023a), 13b CodeLlama and 13b CodeLlama-Instruct (Rozière et al., 2024). Llama-1 has a significantly smaller context window, so it had to read shorter walkthrough prefixes and answer fewer questions. The results of comparing different Llamas are in Figure 8, Figure 9, Table 11, and Table 12. As we can see, both CodeLlama and CodeLlama-Instruct outperform Llama-2, and Llama-1 performs the best in the group. In the main paper, we present the results of Llama-2-base because it is commonly considered to be the standard choice of the Llama series.
(a) On easy (left) and hard (right) DF questions.
(b) On easy (left) and hard (right) RF questions.
Figure 8: Success rates of the Llama-1 and Llama-2 on DF (8(a)) and RF (8(b)) questions, averaged over all $53$ mazes.
(a) On easy (left) and hard (right) DF questions.
(b) On easy (left) and hard (right) RF questions.
Figure 9: Reasoning accuracy of the Llama-1 and Llama-2 on DF (9(a)) and RF (9(b)) questions, averaged over all $53$ mazes.
Method Llama-1 Llama-2 Code-Llama Code-Llama-Instruct $\overline{{\rm HARD}}|$
Llama-1 * 0.29 | 0.25 0.34 | 0.38 0.34 | 0.38 *
Llama-2 0.47 | 0.65 * 0.24 | 0.30 0.24 | 0.30 *
Code-Llama 0.65 | 0.67 0.60 | 0.44 * 0.31 | 0.31 *
Code-Llama-Instruct 0.65 | 0.67 0.60 | 0.44 0.58 | 0.58 * *
|$\underline{{\rm EASY}}$
(a) Pairwise comparison on easy (lower left) and hard (higher right) DF questions.
Method Llama-1 Llama-2 Code-Llama Code-Llama-Instruct $\overline{{\rm HARD}}|$
Llama-1 * 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 *
Llama-2 0.05 | 0.18 * 0.00 | 0.01 0.00 | 0.01 *
Code-Llama 0.03 | 0.20 0.04 | 0.03 * 0.01 | 0.01 *
Code-Llama-Instruct 0.03 | 0.20 0.04 | 0.03 0.03 | 0.03 * *
|$\underline{{\rm EASY}}$
(b) Pairwise comparison on easy (lower left) and hard (higher right) RF questions.
Table 11: Llamas’ success rates on DF and RF questions broken down into pairwise comparison.
Method Llama-1 Llama-2 Code-Llama Code-Llama-Instruct $\overline{{\rm HARD}}|$
Llama-1 * 0.09 | 0.01 0.16 | 0.19 0.16 | 0.19 *
Llama-2 0.25 | 0.40 * 0.02 | 0.06 0.02 | 0.06 *
Code-Llama 0.47 | 0.47 0.34 | 0.18 * 0.07 | 0.07 *
Code-Llama-Instruct 0.47 | 0.47 0.34 | 0.18 0.32 | 0.32 * *
|$\underline{{\rm EASY}}$
(a) Pairwise comparison on easy (lower left) and hard (higher right) DF questions.
Method Llama-1 Llama-2 Code-Llama Code-Llama-Instruct $\overline{{\rm HARD}}|$
Llama-1 * 0.00 | 0.00 0.00 | 0.00 0.00 | 0.00 *
Llama-2 0.05 | 0.18 * 0.00 | 0.00 0.00 | 0.00 *
Code-Llama 0.02 | 0.18 0.03 | 0.03 * 0.00 | 0.00 *
Code-Llama-Instruct 0.02 | 0.18 0.03 | 0.03 0.02 | 0.02 * *
|$\underline{{\rm EASY}}$
(b) Pairwise comparison on easy (lower left) and hard (higher right) RF questions.
Table 12: Llamas’ reasoning accuracies on DF and RF questions broken down into pairwise comparison.
C.6 Human Performance Details
To measure human agreement, we computed the mean square error (MSE) for each metric across the questions that were answered by more than one human rater. As shown in Table 13, the human agreement turns out to be high.
All Easy Hard
Task Type SR RA SR RA SR RA
Route Finding 0.1442 0.3333 0.0425 0.1429 0.5000 1.0000
Destination Finding 0.0000 0.6000 0.0000 0.5714 0.0000 0.6667
Table 13: Human agreement measured by MSE. SR stands for success rate, while RA denotes reasoning accuracy.
C.7 More Results About Playing Minigames
In section 3.5, we have shown that a strong mapping and navigation ability can help an LLM achieve better performance in playing the minigames. Now we show the fine-grained results for that set of experiments. Precisely, Table 14 shows the spelled-out success rates of GPT-3.5 and GPT-4 on each maze, where $M/N$ means that $M$ of $N$ minigames were successfully played by this model. Since GPT-4 by default has a larger context window size (8K) than GPT-3.5 (4K), in order to make a fair comparison, we restrict the context window size to 4K for both GPTs.
Games GPT-3.5 GPT-4
w/o maps w/ maps w/o maps w/ maps
905 0/0 0/0 0/0 0/0
advent 3/9 3/9 4/9 5/9
adventureland 1/7 1/7 3/7 5/7
afflicted 0/12 1/12 4/12 4/12
anchor 0/2 0/2 2/2 2/2
awaken 0/7 0/7 3/7 7/7
balances 0/3 1/3 2/3 2/3
ballyhoo 1/6 1/6 3/6 2/6
curses 2/7 3/7 2/7 6/7
cutthroat 4/5 5/5 4/5 4/5
deephome 3/10 6/10 8/10 10/10
detective 4/6 5/6 6/6 6/6
dragon 0/6 1/6 0/6 2/6
enchanter 0/1 1/1 1/1 1/1
enter 0/0 0/0 0/0 0/0
gold 1/7 2/7 3/7 4/7
hhgg 0/0 0/0 0/0 0/0
hollywood 2/6 2/6 4/6 5/6
huntdark 0/0 0/0 0/0 0/0
infidel 1/4 2/4 2/4 2/4
inhumane 1/2 2/2 2/2 2/2
jewel 0/10 1/10 4/10 7/10
karn 0/8 0/8 4/8 5/8
library 1/5 2/5 5/5 5/5
loose 3/4 3/4 3/4 3/4
lostpig 0/1 0/1 0/1 1/1
ludicorp 1/13 2/13 9/13 10/13
lurking 1/4 0/4 0/4 2/4
moonlit 0/2 2/2 2/2 2/2
murdac 1/2 1/2 2/2 2/2
night 5/16 8/16 9/16 13/16
omniquest 1/11 1/11 1/11 7/11
partyfoul 0/2 1/2 1/2 1/2
pentari 4/13 3/13 4/13 8/13
planetfall 3/6 4/6 3/6 5/6
plundered 0/0 0/0 0/0 0/0
reverb 0/2 0/2 1/2 1/2
seastalker 0/2 0/2 0/2 2/2
sherlock 3/3 3/3 3/3 3/3
snacktime 0/3 0/3 3/3 3/3
sorcerer 4/5 4/5 3/5 4/5
spellbrkr 1/1 1/1 1/1 1/1
spirit 4/14 5/14 8/14 10/14
temple 0/2 1/2 1/2 2/2
trinity 0/0 0/0 0/0 0/0
tryst205 0/3 1/3 0/3 1/3
wishbringer 0/1 1/1 0/1 0/1
yomomma 0/4 0/4 0/4 1/4
zenon 1/5 3/5 5/5 5/5
zork1 2/18 3/18 10/18 13/18
zork2 4/9 5/9 7/9 8/9
zork3 0/6 1/6 0/6 4/6
ztuu 0/9 0/9 7/9 7/9
Table 14: The experiment results (# successful / # answerable minigames) of each model.
Appendix D Future Directions
Our MANGO benchmark sets up a leaderboard, measuring the mapping and navigation abilities of LLMs. An interesting future direction is to investigate how the internal representations of each LLM have—if any—captured the structures of the maps. For example, one may probe the LLM representations and examine whether these representations are predictive of certain characteristics of the maze, similar to the experiments of Li et al. (2023) in board games. This can potentially reveal critical insights into the underlying mechanism of an LLM to represent spatial relations, benefiting their application into real-world scenarios.
The second interesting direction is to investigate how low-cost-adaptation can improve the performance of an LLM. For example, can an LLM, trained on a limited set of mazes, generalize its knowledge and reasoning capability to unseen mazes? Findings in this direction could enhance their usefulness in dynamic environments.
Another direction is to upgrade the MANGO benchmark by enriching its spatial and structural configurations on top of the current maps. These configurations include:
  • spatial notations (e.g., distance in meters, area in square meters) such that one would need complex movements to achieve a target (e.g., not "north" but "north 3 meters");
  • notions of facing directions and rotations such that one would need to turn to switch facing directions.
Such upgrades will enhance its practical application in more complex scenarios. They will also enable investigation in interesting topics, including
  • LLMs’ generalization to significantly out-of-distribution navigation scenarios (e.g., different structures, styles, and configurations);
  • LLMs’ robustness to distracting context.